OpenStudy (anonymous):

simplify into a single fraction. (2x/(x+5)) / (6x^2/2x+10)

OpenStudy (anonymous):

\[\frac{2x}{x+5}\times \frac{2x+10}{6x^2}\] os a start

OpenStudy (anonymous):

then factor to get \[\frac{2x\times 2(x+5)}{(x+5)6x^2}\]

OpenStudy (anonymous):

finally cancel and get \[\frac{2}{3x}\]